OpenStudy (anonymous):

find the point at which the line intersects the plane: x=1+2T y=4T z=2-3T, x+2y-z+1=0

OpenStudy (anonymous):

(1+2T) + 2(4T) - (2-3T) + 1 = 0

OpenStudy (anonymous):

yeah i got that far, but then i get T=0 when i solve

OpenStudy (anonymous):

x = 1 + 2(0) y = 4 (0) z = 2 - 3(0)

OpenStudy (anonymous):

okay got it! thanks!

OpenStudy (anonymous):

np